How to Kiss an Angel

Christoph Benzmüller FR Informatik, Saarland Univ.

chris@ags.uni-sb.de

SEKI Working-Paper SWP–2006–02

SEKI Working-Paperhttp://wirth.bplaced.net/seki.html ISSN 1860-5931

SEKI is published by the following institutions:

German Research Center for Artificial Intelligence (DFKI GmbH), Germany

• Robert Hooke Str. 5, D–28359 Bremen
• Trippstadter Str. 122, D–67663 Kaiserslautern
• Campus D 3 2, D–66123 Saarbrücken

Jacobs University Bremen, School of Engineering & Science, Campus Ring 1, D–28759 Bremen, Germany

Universität des Saarlandes, FR 6.2 Informatik, Campus, D–66123 Saarbrücken, Germany

SEKI Editor:

Claus-Peter Wirth

E-mail: wirth@logic.at
WWW: http://wirth.bplaced.net

Please send surface mail exclusively to:

DFKI Bremen GmbH Safe and Secure Cognitive Systems Cartesium Enrique Schmidt Str. 5 D–28359 Bremen Germany

This SEKI Working-Paper was internally reviewed by:

Erica Melis FR Informatik, Universität des Saarlandes, D–66123 Saarbrücken, Germany

E-mail: melis@activemath.org
WWW: http://www.activemath.org/~melis

How to Kiss an Angel

Christoph Benzmüller FR Informatik, Saarland Univ.

                  chris@ags.uni-sb.de

May 27, 2014

Abstract

We solve the problem of kissing an angel effectively and more efficiently than voodoo man. Some reading of Immanuël Kant, however, is required as a precondition.

Keywords: German Idealism, Dyslexia, Kissing, Angels, Angles, Kant, MSC 03D20, MSC 68Q42, MSC 16S15

1 Introduction and Conclusion

Kiss an angle and get dyslexia!